a sample of 600 respondents the monthly income of the respondents follow a normal distribution with its mean and standard deviation at 30,000 and 6000 respectively
What is the probability that monthly income is between 20000 and 34000
Probability that a monthly income is more than 32,000
How many people Expected to have the income between 20,000 and 34000
What is the probability that the monthly income is less than 24,000
1
Expert's answer
2021-07-26T14:27:20-0400

"N=600 \\

\mu=30000 \\

\sigma = 6000"

Is it justifiable to exclude homemakers’ value creation from national income?

"P(20000<X<34000) = P(X<34000) -P(X<20000) \\

= P(Z< \frac{34000-30000}{6000} -P(Z< \frac{20000-30000}{6000}) \\

=P(Z< 0.666) -P(Z< -1.666) \\

= 0.7476 -0.0478 \\

= 0.6998"

Probability that a monthly income is more than 32,000

"P(X>32000) = 1 -P(X<32000) \\

= 1 -P(Z< \frac{32000-30000}{6000}) \\

= 1 -P(Z< 0.3333) \\

= 1 -06304 \\

= 0.3696"

How many people Expected to have the income between 20,000 and 34000

Number of people "= N \times P(20000<X<34000)"

"= 600 \times 0.6998 \\

= 419.88 ≈420"

What is the probability that the monthly income is less than 24,000

"P(X< 24000) = P(Z< \frac{24000-30000}{6000}) \\

= P(Z< -1) \\

= 0.1586"